# Scalable hybrid quantum Monte Carlo simulation of U(1) gauge field coupled to fermions on GPU

### v4

@@ I Results
Figure 2: The speedup achieved by customized CUDA kernel and CUDA Graph, and the computational complexity of HQMC. Panel (a) shows the speedup achieved by customized CUDA kernel for different system sizes $L$, with $V_{s}=L^{2}$ and $N_{\tau}=10L$. The blue axis on the left represents the latency of the Monte Carlo sweep which is average time cost per generated sample. The light blue line represents the baseline performance, where the PyTorch sparse-matrix-vector multiplication is applied to compute $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$. The dark blue line is the latency of the customized CUDA kernel, where the computations $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$ are replaced with optimized CUDA kernel. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size. (b) shows the speedup achieved by CUDA Graph, for different system sizes $L$. The light blue line represents the baseline performance with CUDA Graph turned off, while the dark blue line is the latency when CUDA Graph is turned on. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size. (c) shows the scaling behavior of the HQMC sample generation latency and HQMC measurement latency using stochastic estimation (SE), as functions of $N_{\tau}V_{s}\propto L^{3}$, where $N_{\tau}=\beta/\Delta_{\tau}=10L$. The latency (second per sample) is defined as the time spent to generate a gauge sample, or the time to compute a fermionic observable from a gauge sample for SE (the SE latency here is tested on spin-spin correlation using $N_{\textmd{rv}}=40$). The blue and orange line are the linear fitting of five large-size points. They show that the HQMC algorithm has a time complexity of $O((L^{3})^{0.9})$, while the stochastic estimation has a time complexity of $O((L^{3})^{1.3})$. As a contrast, the DQMC has a time complexity of $O(N_{\tau}V_{s}^{3})=O(L^{7})$, which is plotted as the green dashed line. The starting point of the green dashed line is $0.38$ second per sweep at $L^{3}=10^{3}$, obtained from DQMC benchmark.
@@
$C_{\textmd{S}}^{\alpha\beta}(i,j)$ $=\langle S^{\alpha\beta}(i)S^{\beta\alpha}(j)\rangle$ (4)
@@
Other than the gauge invariant fermionic observables, we also measure the correlation of the magnetic flux operator $\theta_{\square}:=\textmd{curl}\phi=\sum_{ij\in\textmd{edge}(\Box)}\phi_{ij}$, where $\phi_{ij}=-\phi{ji}$ and the summation is along the positive directions shown in Fig. 1. We focus on the imaginary time dynamics. Since, in the static limit ($J=0$) of the QED3, the ground energy magnetic configuration is known to be $\pi$-flux on a square lattice [41]. When $J$ is turned on, the fluctuation of magnetic flux is centered also around $\pi$-flux. So, to study the correlation of the magnetic fluxes, we measure the correlation of its sine value instead, in order to have a clean background, i.e. $\langle\sin\theta_{\square}\rangle=0$. We define
@@
Figure 3: The correlation functions for non-compact QED3 with $K=1$ and $J=1.25$. Panel (a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$, for various cubic lattice sizes shown as a format of $N_{\tau}\times L^{2}$ in the legends in both panel (a) and (b). Notice that $C_{S}^{\uparrow\downarrow}=C_{S}^{\downarrow\uparrow}$, so we only need to measure $C_{S}^{\uparrow\downarrow}$. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray data points show the results computed from the DQMC. The inset plots the zoomed in view of the dashed region, which shows that DQMC results agrees with the HQMC result within the error bars. The HQMC data points at the largest lattice size, above the noise level of $10^{-6}$, are manually fit by the black solid line, which displays a power law decay as $1/r^{3.8}$. Panel (b) shows the log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$ for various lattice sizes. The data points of the largest lattice size, before entering the noise level at $10^{-5}$, are manually fit by the black solid line, and also displays a power law decay with $1/r^{3.7}$. Panel (c) shows the log plot of the flux-flux correlation as a function of imaginary time $\tau$. At the large $\tau$ and the large system sizes, the black solid line shows the manual fitting which displays the clear linear behavior indicating a power-law decay as $\tau^{-4.0}$. As a comparison, the black dashed line shows the $\tau^{-3.0}$ power law behavior. The inset plots zoomed in view of the data in the dashed squared region. Panels (d-f) show the operator scaling exponent $\Delta_{S}$, $\Delta_{B}$, $\Delta_{\textmd{flux}}$ for different lattice sizes obtained from the linear fitting of the data points shown in panels (a-c), respectively (the fit lines are not shown). To obtain the exponent shown in panels (d, e), the data points in panels (a, b) above the noise level for each lattice size are linearly fit, and the slope gives the scaling exponent shown as the blue dots in panels (d, e), which are then extrapolated to $1/L\to 0$, leveraging the rational curves $y=y_{0}+\frac{b\cdot x^{d}}{1+c\cdot x}$ (red dashed curves), where $y_{0},b,c,d$ are free parameters. Similarly, the blue dots in panel (f) are the exponents obtained from the linear fitting of the data points in panel (c) for each lattice size. The data points used in the linear fitting are the linear region above the noise level, which are taken from the $C_{\textmd{flux}}$ interval of $10^{-3},2\times 10^{-2}]$. Again, the red dashed curve and the dot show the $1/L\to 0$ extrapolation.
@@ III Methods
$M_{\alpha}$ $=\begin{pmatrix}\mathbf{1}&0&0&\cdots&0&B_{N_{\tau},\alpha}\\ -B_{1,\alpha}&\mathbf{1}&0&\cdots&0&0\\ 0&-B_{2,\alpha}&\mathbf{1}&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\cdots&\mathbf{1}&0\\ 0&0&0&\cdots&-B_{N_{\tau}-1,\alpha}&\mathbf{1}\end{pmatrix}$. (8)
@@
    $\dot{p}_{ij,\tau}$ $=-\frac{\partial H}{\partial\phi_{ij,\tau}}$ (13)
@@
$p_{n-\frac{1}{2}}$ $=p_{n-1}+F\left(\phi_{n-1}\right)\frac{\Delta t}{2}$
$\phi_{n}$ $=\phi_{n-1}+2p_{n-\frac{1}{2}}\Delta t$ (16)
@@
Here, the force $F(\phi):=-\frac{\partial H}{\partial\phi}$ consists of both the bosonic force $\frac{\partial S_{B}}{\partial\phi}$ and the fermionic force $\frac{\partial}{\partial\phi}\left\{\eta^{\dagger}\left[(M^{\dagger}(\phi)M(\phi)\right]^{-1}\eta\right\}$, the derivation of which is shown in SM Sec. SI. Calculation of Fermionic Force for QED3. In this calculation, generically a matrix vector of the form $\left[M^{\dagger}M\right]^{-1}v$ is involved, which is equivalent to solving a linear system $\left[M^{\dagger}M\right]X=v$, and will be calculated using a preconditioned conjugate gradient (pCG) algorithm, which will be explained in section CG Method and Preconditioner below. In this Leapfrog method, the time step size $\Delta t$ and the number of leapfrog steps $N_{\textmd{leapfrog}}$ in a Monte Carlo sweep are two tunable parameters. Larger $N_{\textmd{leapfrog}}$ benefits from larger change in the proposed configuration, but suffers from the larger accumulated error in the iterative process. We choose $N_{\textmd{leapfrog}}=3$ in our implementation. The time step size $\Delta t$, on the other hand, is adaptively tuned to maintain a moderate acceptance rate.
@@
$C(\Delta)=\frac{1}{N}\sum_{i}C(i,i+\Delta)$ (17)
@@
$X_{n+1}$ $=X_{n}+\zeta_{n}d_{n}$
$r_{n+1}$ $=r_{n}-\zeta_{n}Od_{n}$
$\rho_{n}$ $=\left\langle r_{n+1},\tilde{O}^{-1}r_{n+1}\right\rangle/\left\langle r_{n},\tilde{O}^{-1}r_{n}\right\rangle$ (20)
$d_{n+1}$ $=\tilde{O}^{-1}r_{n+1}+\rho_{n}d_{n}$
@@ III.0.1 Matrix-free representation of $M^{\dagger}Mv$ multiplication
$e^{V_{\textmd{fam}_{n}}}$ $=\bigoplus_{ij\in\textmd{fam}_{n}}e^{\Delta_{\tau}T_{ij}}$ (24)
$=\bigoplus_{ij\in\textmd{fam}_{n}}\left(\mathbf{1}\cosh\Delta_{\tau}+T_{ij}\sinh\Delta_{\tau}\right)$ (25)
@@ Acknowledgments
Scalable Hybrid quantum Monte Carlo simulation of QED3 coupled to fermions on GPU
In this supplemental material, we provide the technical details concerning the QMC measurements of conserved currents (Sec. SII) and the finite size scaling of the order parameters at the easy-plane DQCP (Sec. SIII), the derivation of the fermionic correlation functions (Sec. SIV), and the analysis of autocorrelation time (Sec. SV).
@@
Figure S5: The correlation functions for compact QED3 with $K=1$ and $J=1.25$. Panel (a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$, for various cubic lattice sizes shown in legends in panel (a) and (b) as the format of $N_{\tau}\times L^{2}$. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray data points show the results computed from the DQMC. The inset plots the zoomed in view of the dashed region, which shows that DQMC results agrees with the HQMC result within the error bars. The HQMC data of largest lattice size, before entering the noise level of $10^{-6}$, are manually fitted by the black line, which displays a power law decay with scaling dimension of around $3.8$. Panel (b) shows the log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$. The data points of largest lattice size, above the noise level of $10^{-5}$, are manually fitted by the black solid line and displays a power law decay as $1/r^{3.7}$. Panel (c) shows the log plot of the flux-flux correlation as a function of imaginary time $\tau$. At the large $\tau$ and the large system sizes, the black solid line shows the manual fitting which displays a power law decay as $\tau^{-4.1}$. As a comparison, the black dashed line shows the $\tau^{-3.0}$ power law behavior. The inset plots zoomed in view of the data in the dashed squared region. Panels (d-f) show the operator scaling exponent $\Delta_{S}$, $\Delta_{B}$, $\Delta_{\textmd{flux}}$ for different lattice sizes obtained from the linear fitting of the data shown in panels (a-c) respectively (the fit lines are not shown). To obtain the exponent shown in panels (d, e), in panels (a, b), the data points above the noise level for each lattice sizes are linearly fit, and the slope gives the scaling exponent shown as the blue dots in panel (d, e), which are then extrapolated to $1/L\to 0$ shown as the red dashed curve and the dot. Similarly, the blue dots in panel (f) are the exponents obtained from the linear fitting of the data points in panel (c) for each lattice size. The data points used in the linear fitting are the linear region above the noise level, which are taken from the $C_{\textmd{flux}}$ interval of $10^{-3},2\times 10^{-2}]$. Again, the red dashed curve and the dot show the $1/L\to 0$ extrapolation.

### v5

@@ I Results
Figure 2: The speedup achieved by customized CUDA kernel and CUDA Graph, and the computational complexity of HQMC. Panel (a) shows the speedup achieved by customized CUDA kernel for different system sizes $L$, with $V_{s}=L^{2}$ and $N_{\tau}=10L$. The blue axis on the left represents the latency of the Monte Carlo sweep which is average time cost per generated sample. The light blue line represents the baseline performance, where the PyTorch sparse-matrix-vector multiplication is applied to compute $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$. The dark blue line is the latency of the customized CUDA kernel, where the computations $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$ are replaced with optimized CUDA kernel. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size. Panel (b) shows the speedup achieved by CUDA Graph, for different system sizes $L$. The light blue line represents the baseline performance with CUDA Graph turned off, while the dark blue line is the latency when CUDA Graph is turned on. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size. Panel a(c) shows the scaling behavior of the HQMC sample generation latency and HQMC measurement latency using stochastic estimation (SE), as functions of $N_{\tau}V_{s}\propto L^{3}$, where $N_{\tau}=\beta/\Delta_{\tau}=10L$. The latency (second per sample) is defined as the time spent to generate a gauge sample, or the time to compute a fermionic observable from a gauge sample for SE (the SE latency here is tested on spin-spin correlation using $N_{\textmd{rv}}=40$). The blue and orange line are the linear fitting of five large-size points. They show that the HQMC algorithm has a time complexity of $O((L^{3})^{0.9})$, while the stochastic estimation has a time complexity of $O((L^{3})^{1.3})$. As a contrast, the DQMC has a time complexity of $O(N_{\tau}V_{s}^{3})=O(L^{7})$, which is plotted as the green dashed line. The starting point of the green dashed line is $0.38$ second per sweep at $L^{3}=10^{3}$, obtained from DQMC benchmark.
@@
$C_{\textmd{S}}^{\alpha\beta}(i,j)$ $=\langle S^{\alpha\beta}(i)S^{\beta\alpha}(j)\rangle,$ (4)
@@
Other than the gauge invariant fermionic observables, we also measure the correlation of the magnetic flux operator $\theta_{\square}:=\textmd{curl}\phi=\sum_{ij\in\textmd{edge}(\Box)}\phi_{ij}$, where $\phi_{ij}=-\phi_{ji}$ and the summation is along the positive directions shown in Fig. 1. We focus on the imaginary time dynamics. Since, in the static limit ($J=0$) of the QED3, the ground energy magnetic configuration is known to be $\pi$-flux on a square lattice [40]. When $J$ is turned on, the fluctuation of magnetic flux is centered also around $\pi$-flux. So, to study the correlation of the magnetic fluxes, we measure the correlation of its sine value instead, in order to have a clean background, i.e. $\langle\sin\theta_{\square}\rangle=0$. We define
@@
Figure 3: The correlation functions for non-compact QED3 with $K=1$ and $J=1.25$. Panel (a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$, for various cubic lattice sizes shown as a format of $N_{\tau}\times L^{2}$ in the legends in both panel (a) and (b). Notice that $C_{S}^{\uparrow\downarrow}=C_{S}^{\downarrow\uparrow}$, so we only need to measure $C_{S}^{\uparrow\downarrow}$. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray data points show the results computed from the DQMC. The inset plots the zoomed in view of the dashed region, which shows that DQMC results agrees with the HQMC result within the error bars. The HQMC data points at the largest lattice size, above the noise level of $10^{-6}$, are manually fit by the black solid line, which displays a power law decay as $1/r^{3.8}$. Panel (b) shows the log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$ for various lattice sizes. The data points of the largest lattice size, before entering the noise level at $10^{-5}$, are manually fit by the black solid line, and also displays a power law decay with $1/r^{3.7}$. Panel (c) shows the log plot of the flux-flux correlation as a function of imaginary time $\tau$. At the large $\tau$ and the large system sizes, the black solid line shows the manual fitting which displays the clear linear behavior indicating a power-law decay as $\tau^{-4.0}$. As a comparison, the black dashed line shows the $\tau^{-3.0}$ power law behavior. The inset plots zoomed in view of the data in the dashed squared region. Panels (d-f) show the operator scaling exponent $\Delta_{S}$, $\Delta_{B}$, $\Delta_{\textmd{flux}}$ for different lattice sizes obtained from the linear fitting of the data points shown in panels (a-c), respectively (the fit lines are not shown). To obtain the exponent shown in panels (d, e), the data points in panels (a, b) above the noise level for each lattice size are linearly fit, and the slope gives the scaling exponent shown as the blue dots in panels (d, e), which are then extrapolated to $1/L\to 0$, leveraging the rational curves $y=y_{0}+\frac{b\cdot x^{d}}{1+c\cdot x}$ (red dashed curves), where $y_{0},b,c,d$ are free parameters. Similarly, the blue dots in panel (f) are the exponents obtained from the linear fitting of the data points in panel (c) for each lattice size. The data points used in the linear fitting are the linear region above the noise level, which are taken from the $C_{\textmd{flux}}$ interval of $1\times 10^{-3},2\times 10^{-2}]$. Again, the red dashed curve and the dot show the $1/L\to 0$ extrapolation.
@@ III Methods
$M_{\alpha}$ $=\begin{pmatrix}\mathbf{1}&0&0&\cdots&0&B_{N_{\tau},\alpha}\\ -B_{1,\alpha}&\mathbf{1}&0&\cdots&0&0\\ 0&-B_{2,\alpha}&\mathbf{1}&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\cdots&\mathbf{1}&0\\ 0&0&0&\cdots&-B_{N_{\tau}-1,\alpha}&\mathbf{1}\end{pmatrix}$, (8)
@@
    $\dot{p}_{ij,\tau}$ $=-\frac{\partial H}{\partial\phi_{ij,\tau}},$ (13)
@@
$p_{n-\frac{1}{2}}$ $=p_{n-1}+F\left(\phi_{n-1}\right)\frac{\Delta t}{2},$
$\phi_{n}$ $=\phi_{n-1}+2p_{n-\frac{1}{2}}\Delta t,$ (16)
@@
Here, the force $F(\phi):=-\frac{\partial H}{\partial\phi}$ consists of both the bosonic force $\frac{\partial S_{B}}{\partial\phi}$ and the fermionic force $\frac{\partial}{\partial\phi}\left\{\eta^{\dagger}\left[(M^{\dagger}(\phi)M(\phi)\right]^{-1}\eta\right\}$, the derivation of which is shown in Sec. SI of SM [2]. In this calculation, generically a matrix vector of the form $\left[M^{\dagger}M\right]^{-1}v$ is involved, which is equivalent to solving a linear system $\left[M^{\dagger}M\right]X=v$, and will be calculated using a preconditioned conjugate gradient (pCG) algorithm, which will be explained in section CG Method and Preconditioner below. In this Leapfrog method, the time step size $\Delta t$ and the number of leapfrog steps $N_{\textmd{leapfrog}}$ in a Monte Carlo sweep are two tunable parameters. Larger $N_{\textmd{leapfrog}}$ benefits from larger change in the proposed configuration, but suffers from the larger accumulated error in the iterative process. We choose $N_{\textmd{leapfrog}}=3$ in our implementation. The time step size $\Delta t$, on the other hand, is adaptively tuned to maintain a moderate acceptance rate.
@@
$C(\Delta)=\frac{1}{N}\sum_{i}C(i,i+\Delta),$ (17)
@@
$X_{n+1}$ $=X_{n}+\zeta_{n}d_{n},$
$r_{n+1}$ $=r_{n}-\zeta_{n}Od_{n},$
$\rho_{n}$ $=\left\langle r_{n+1},\tilde{O}^{-1}r_{n+1}\right\rangle/\left\langle r_{n},\tilde{O}^{-1}r_{n}\right\rangle,$ (20)
$d_{n+1}$ $=\tilde{O}^{-1}r_{n+1}+\rho_{n}d_{n},$
@@ III.0.1 Matrix-free representation of $M^{\dagger}Mv$ multiplication
$e^{V_{\textmd{fam}_{n}}}$ $=\bigoplus_{ij\in\textmd{fam}_{n}}e^{\Delta_{\tau}T_{ij}},$ (24)
$=\bigoplus_{ij\in\textmd{fam}_{n}}\left(\mathbf{1}\cosh\Delta_{\tau}+T_{ij}\sinh\Delta_{\tau}\right),$ (25)
@@ Acknowledgments
In this supplemental material, we provide the technical details concerning the Calculation of Fermionic Force for QED3 (Sec. SI), the QMC measurements of conserved currents (Sec. SII), the finite size scaling of the order parameters at the easy-plane DQCP (Sec. SIII), the derivation of the fermionic correlation functions (Sec. SIV), and the analysis of autocorrelation time (Sec. SV).
@@
Figure S5: The correlation functions for compact QED3 with $K=1$ and $J=1.25$. Panel (a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$, for various cubic lattice sizes shown in legends in panel (a) and (b) as the format of $N_{\tau}\times L^{2}$. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray data points show the results computed from the DQMC. The inset plots the zoomed in view of the dashed region, which shows that DQMC results agrees with the HQMC result within the error bars. The HQMC data of largest lattice size, before entering the noise level of $10^{-6}$, are manually fitted by the black line, which displays a power law decay with scaling dimension of around $3.8$. Panel (b) shows the log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$. The data points of largest lattice size, above the noise level of $10^{-5}$, are manually fitted by the black solid line and displays a power law decay as $1/r^{3.7}$. Panel (c) shows the log plot of the flux-flux correlation as a function of imaginary time $\tau$. At the large $\tau$ and the large system sizes, the black solid line shows the manual fitting which displays a power law decay as $\tau^{-4.1}$. As a comparison, the black dashed line shows the $\tau^{-3.0}$ power law behavior. The inset plots zoomed in view of the data in the dashed squared region. Panels (d-f) show the operator scaling exponent $\Delta_{S}$, $\Delta_{B}$, $\Delta_{\textmd{flux}}$ for different lattice sizes obtained from the linear fitting of the data shown in panels (a-c) respectively (the fit lines are not shown). To obtain the exponent shown in panels (d, e), in panels (a, b), the data points above the noise level for each lattice sizes are linearly fit, and the slope gives the scaling exponent shown as the blue dots in panel (d, e), which are then extrapolated to $1/L\to 0$ shown as the red dashed curve and the dot. Similarly, the blue dots in panel (f) are the exponents obtained from the linear fitting of the data points in panel (c) for each lattice size. The data points used in the linear fitting are the linear region above the noise level, which are taken from the $C_{\textmd{flux}}$ interval of $1\times 10^{-3},2\times 10^{-2}]$. Again, the red dashed curve and the dot show the $1/L\to 0$ extrapolation.